OpenStudy (hothotleg):

Fan/Medal/Testimony Marlena was asked to find an expression that is not equivalent to 2 12. Which of the following is not equivalent to the given expression? a. (2^2)^6 b. (2^8)^4 c.(2^6)(2^6) d.(2^3)(2^9)

OpenStudy (mjdennis):

Can you clarify (clear up) the question? Is that the number 212? Or 2^12?

OpenStudy (hothotleg):

Oh! yes, it is 2^12

OpenStudy (michele_laino):

hint: the subsequent identities hold: \[\huge \begin{gathered} {\left( {{2^a}} \right)^b} = {2^{\left( {a \times b} \right)}} \hfill \\ \hfill \\ {2^a} \times {2^b} = {2^{\left( {a + b} \right)}} \hfill \\ \end{gathered} \]

OpenStudy (hothotleg):

I don't quite understand...But I'll try to figure it out!